### v1

# Climplicit: Climatic Implicit Embeddings for Global Ecological Tasks

Abstract
Deep learning on climatic data holds potential for macroecological applications. However, its adoption remains limited among scientists outside the deep learning community due to storage, compute, and technical expertise barriers. To address this, we introduce Climplicit, a spatio-temporal geolocation encoder pretrained to generate implicit climatic representations anywhere on Earth. By bypassing the need to download raw climatic rasters and train feature extractors, our model uses ×$\times 1000$ fewer disk space and significantly reduces computational needs for downstream tasks. We evaluate our Climplicit embeddings on biomes classification, species distribution modeling, and plant trait regression. We find that linear probing our Climplicit embeddings consistently performs better or on par with training a model from scratch on downstream tasks and overall better than alternative geolocation encoding models.
1 Introduction
Ecology studies how species interact with their environment, and climate is a key factor in shaping habitats and distributions (Karger et al., 2017). Recently, deep learning has been successfully applied to the modeling of species distribution from citizen science data (Cole et al., 2023; Botella et al., 2023; Brun et al., 2024; Dollinger et al., 2024). These models rely on climatic variables, which are derived from vast amounts of raw weather data collected by global stations and aggregated into climatic reanalysis datasets. Although deep learning is well suited to leverage such data, its adoption in ecology remains limited due to high computational, storage, and expertise requirements.
To this end, several works focus on pretraining lightweight, off-the-shelf geolocation encoders for downstream tasks (Rolf et al., 2021; Mai et al., 2023a; Klemmer et al., 2023; Vivanco Cepeda et al., 2024; Agarwal et al., 2024; Sastry et al., 2025). However, existing models often suffer from low resolution or produce inconsistent embeddings (see Appendix F.3, Figure 5), limiting their effectiveness. To address these shortcomings, we introduce Climplicit, a spatio-temporal location encoder that generates implicit climatic representations for any land location and month. By pretraining on CHELSA climatologies (Karger et al., 2017; 2018), our model captures fine-grained environmental patterns with greater consistency and resolution. Our Climplicit backbone is based on residual SIREN (ReSIREN) blocks, a novel extension of SIREN (Sitzmann et al., 2020) with residual connections that improve scalability. Although previous works have explored compressing climatic data into neural networks (Huang and Hoefler, 2022; Naour et al., 2023; Liu et al., 2023; Xu et al., 2024), our approach specifically focuses on a high-resolution, lightweight location encoder. We validate Climplicit embeddings on biodiversity-related tasks and provide ablation studies for our key design choices. Code is available at https://github.com/ecovision-uzh/climplicit.
2 Methodology
Let $i\in\mathbb{N}$ be an index over discretized month-locations on Earth. We define a location encoder as a neural network LE that transforms the positional encoding $p_{i}$ of a month-coordinate triplet $(\lambda_{i},\phi_{i},m_{i})$ with longitude $\lambda_{i}\in[-1,1]$, latitude $\phi_{i}\in[-1,1]$ and month $m_{i}\in\{1,...,12\}$ into an embedding $e_{i}=\text{{LE}}(p_{i})$.
2.1 Direct Positional Embedding with H-SIREN
Our positional encoding extends with a temporal dimension the Direct embedding defined by Rußwurm et al. (2024):
$p_{i}=\Big{[}\>\lambda_{i},\>\phi_{i},\>\text{sin}(2\pi*m_{i}/12),\>\text{cos}(2\pi*m_{i}/12)\>\Big{]}\;\in\;[-1,1]^{4}\;.$
We choose the Direct embedding because of its low computational cost and sufficient representational power in conjunction with the SIREN network (Rußwurm et al., 2024). It is also free of hyperparameters that predefine frequencies, unlike Fourier Features (Tancik et al., 2020), Spherical Harmonics (Rußwurm et al., 2024) and Sphere2Vec (Mai et al., 2023b). The sine and cosine transformations of the months ensure that the values are in $[-1,1]$ and that January and December have close representations. To increase the frequencies accessible to the network, we use the H-SIREN (Gao and Jaiman, 2024) activation function in the first layer (see Appendix B.1).
Figure 1: Left: Contrary to “classic” residual blocks, our ReSIREN residual connection is placed before the non-linearity rather than after. Right: Scaling comparison of SIREN and ReSIREN on the biome classification task (Section D). ReSIREN underperforms at small depths but scales better. See Appendix F.2 for further details.
2.2 Residual Connection for SIREN
Residual connections have become a staple of neural networks since they enabled the first deep architectures (He et al., 2016a). We introduce ReSIREN, a residual extension of SIREN (Sitzmann et al., 2020) that improves scalability and enhances information exchange between layers (He et al., 2016b). Let $L_{j}$, where $j\in\{1,...,D\}$, be the $j$-th layer in a network of depth $D$, consisting of an affine transformation $F_{j}$ and activation function $A_{j}$. Given input $z_{j}$, we define the hidden pre-activation vector as $h_{j}=F_{j}(z_{j})$ and the output as $z_{j+1}=L_{j}(z_{j})$. We introduce residual connections between the hidden vectors:
$h^{\prime}_{j}=\frac{h_{j}+h^{\prime}_{j-1}}{2},\quad z_{j+1}=A_{j}(h^{\prime}_{j}).$
This preserves the normal distribution $\mathcal{N}(0,1)$ of $h_{j}$ required for the stability of SIREN (Sitzmann et al., 2020). Figure 1 illustrates ReSIREN and its improved scaling properties compared to SIREN. See Appendix B.2 and F.2 for more details.
2.3 Climatic Pretraining
[…]
3 Experiments
We assess Climplicit embeddings with linear probing on three macroecological downstream tasks: Biomes classification, species distribution modeling (SDM) and Plant traits regression (see Appendix D for more details). To illustrate the storage- and compute-intensive nature of training that Climplicit aims to mitigate, we train a model from scratch (FS) with a location encoder (Loc) or a CHELSA encoder (CH) for each downstream task. For Loc, we use a 16-layer ReSIREN with the same hyperparameters as Climplicit, but with $[\lambda,\phi]$ as input. For CH, we incorporate CHELSA values from March, June, September, and December, processing them with a 4-layer residual network following Cole et al. (2023). For Loc $+$ CH, we concatenate both embeddings before the probe. Additionally, we compare Climplicit embeddings with those from SatCLIP (Klemmer et al., 2023), Taxabind (Sastry et al., 2025), SINR (Cole et al., 2023), CSP (Mai et al., 2023a), and GeoCLIP (Vivanco Cepeda et al., 2024), which implicitly represent modalities such as satellite imagery, sound, bioclimatic variables, and species occurrence data.
4 Results
Table 1: Linear probing on diverse geolocation representations. Mean and standard deviation of ten random initializations of the linear probe. Best. Second best.
Model Biomes (F1 $\uparrow$) SDM (Acc $\uparrow$) Plant traits (R² $\uparrow$)
FS Loc 73.9 $\pm$ 2.4 2.0 $\pm$ 0.4 42.2 $\pm$ 0.0
FS CH 71.8 $\pm$ 1.9 2.5 $\pm$ 0.1 60.0 $\pm$ 0.3
FS Loc $+$ CH 79.6 $\pm$ 1.7 2.5 $\pm$ 0.1 64.8 $\pm$ 0.4
SatCLIP 68.3 $\pm$ 0.4 1.3 $\pm$ 0.1 61.6 $\pm$ 0.1
Taxabind 59.3 $\pm$ 0.1 3.1 $\pm$ 0.0 56.9 $\pm$ 0.0
SINR 63.1 $\pm$ 0.3 1.7 $\pm$ 0.0 63.5 $\pm$ 0.1
CSP 58.6 $\pm$ 0.4 1.6 $\pm$ 0.1 49.7 $\pm$ 0.3
GeoCLIP 62.7 $\pm$ 0.1 3.5 $\pm$ 0.0 57.9 $\pm$ 0.1
Climplicit (Ours) 78.4 $\pm$ 0.3 3.2 $\pm$ 0.0 70.0 $\pm$ 0.1
We summarize our results in Table 1.
Climplicit representations surpass all FS ones, except for FS Loc $+$ CH on Biomes, which slightly outperforms Climplicit but with a larger confidence interval, highlighting the high variance of representations learned from scratch on small datasets compared to pretrained representations. These results validate the potential of Climplicit for scenarios with minimal storage, compute, or deep learning knowledge. Furthermore, we observe that other geolocation encoders fall short of Climplicit on all tasks, with the exception of GeoCLIP on SDM, which we attribute to the high resolution of GeoCLIP embeddings. This highlights the relevance of high-resolution climatic representations for the downstream tasks at hand. Further quantitative and qualitative results can be found in Appendix F.1, F.2, and F.3.
5 Ablation Studies
We analyze the impact of key design choices in Table 2. SIREN removes the residual connections from our ReSIREN blocks, consistently resulting in a performance drop, with $-1.2$ R2 for plant traits regression. Appendix F.2 further highlights the impact of residual connections with increasing model depth. To explore the impact of learning temporal climatic data representations, All- Months is pretrained to reconstruct the concatenation of all monthly CHELSA variables at each location, while March-Only is pretrained exclusively on March data ($m_{i}=3$). Our results support support the choice of making Climplicit a spatio-temporal location encoder rather than encoding only a single month or the entirety of yearly climatic variables. In No H-SIREN we remove the H-SIREN activation function and observe a performance drop on two of our tasks. rec-CHELSA runs linear probing on the reconstructed CHELSA variables instead of the embeddings before the projection layer. As expected, the latent representation better generalizes compared to reconstructed climatic variables. In CHELSA-CLIP, we pretrain our model with a contrastive objective similar to Klemmer et al. (2023) which proves to produce poorer representations compared to directly regressing CHELSA. Whether this approach extends to more complicated regressions or auto-encoding settings for multi-pixel images is left to future work. Finally, ERA5 compares representations learned from monthly ERA5 (Dee et al., 2011) instead of CHELSA. We find that the ×$\times 30$ higher resolution of CHELSA ($30$ versus $900$ arcsec) allows for learning richer representations for our marcoecological tasks.
Table 2: Ablation analysis with linear probing. Mean and standard deviation of ten random initializations of the linear probe. Best.
Model Biomes (F1 $\uparrow$) SDM (Acc $\uparrow$) Plant traits (R² $\uparrow$)
Climplicit 78.4 $\pm$ 0.3 3.2 $\pm$ 0.0 70.0 $\pm$ 0.1
SIREN 77.5 $\pm$ 0.2 3.1 $\pm$ 0.0 68.8 $\pm$ 0.2
Concat Months 75.9 $\pm$ 0.3 2.6 $\pm$ 0.0 66.0 $\pm$ 0.1
March-Only 78.2 $\pm$ 0.2 2.9 $\pm$ 0.0 62.8 $\pm$ 0.1
No H-SIREN 77.9 $\pm$ 0.2 3.6 $\pm$ 0.0 69.1 $\pm$ 0.1
rec-CHELSA 61.5 $\pm$ 0.2 1.5 $\pm$ 0.0 55.4 $\pm$ 0.1
CH-CLIP 76.5 $\pm$ 0.6 2.3 $\pm$ 0.1 66.9 $\pm$ 0.4
ERA5 63.7 $\pm$ 0.5 1.9 $\pm$ 0.1 68.6 $\pm$ 0.2
6 Discussion and Conclusion
We introduced Climplicit, a spatiotemporal encoder that produces implicit climatic representations anywhere on Earth, aimed at lowering the storage, compute, and machine learning expertise requirements for downstream practitioners. Our model hinges on ReSIREN, a residual extension of SIREN that allows to learn high-resolution representations. Our Climplicit model occupies $99.9\%$ fewer memory than the CHELSA dataset it was trained on, while producing representations that perform on par or better than from-scratch training on several macroecological tasks. Potential limitations of Climplicit embeddings are their implicit nature and lower resolution, compared to the original climatic raster dataset. Finally, in Appendix E, we outline how Climplicit could be extended to provide embeddings for future climatologies under different climate change scenarios, in turn allowing future predictions for the downstream tasks.
Appendix A Impact and Ethical Statement
This work is part of an ongoing project co-designed with ecologists to make deep learning methods more easily accessible to practitioners. Future work will apply the model to more extensive ecological problems. After finishing pretraining, this model has small requirements in regards to compute, storage capacity and know-how, which increases ease of use for non-deep learning experts and reduces the carbon footprint of downstream applications. No GenAI such as ChatGPT was used in the making of this work and writing of this publication.
Appendix B ReSIREN Details
B.1 ReSIREN Activations
Let $L_{j},j\in\{1,...,D\}$ be the j-th layer of the network of depth $D$, consisting of an affine transformation $F_{j}$ and activation function $A_{j}$. The activation functions across the network then are:
$A_{j}(x)=\begin{cases}\text{sin}(\omega_{0}\text{sinh}(2x))&\text{for j = 1}\\ \text{sin}(\omega_{0}x)&\text{for 1 < j < D}\\ \text{Identity}(x),&\text{for j = D}\end{cases}$
Following the results of Sitzmann et al. (2020), we set $\omega_{0}=30$. According to the authors, $\omega_{0}$ increases the input range across multiple periods of sin, which allows the network to converge more quickly and access more frequencies. The special case in the first layer is the H-SIREN function of Gao and Jaiman (2024).
B.2 Distributional Stability of ReSIREN
Figure 2: Experimental probability density function of adding up two independent $\text{Arcsin}(0,1)$. The sum is a ill-behaved distribution that can not be trivially normalized back to $\text{Arcsin}(0,1)$. Therefore, we choose to apply residual connections between to the well-behaved Gaussians in the middle of each SIREN layer.
Let $z_{j}$ be the input to $L_{j}$, $h_{j}=F_{j}(z_{j})$ the hidden pre-activation vector and $z_{j+1}=L_{j}(z_{j})$ the output of the layer. According to Sitzmann et al. (2020), the distribution of $z_{j}$ is $\text{Arcsin}(-1,1)$ and $\mathcal{N}(0,1)$ for $h_{j}$. These distributional properties are essential to the functioning of the network and are enabled by a specific initialization of the weights of $F_{j}$. Adding residual connections the classic way by adding the output of the previously layer to the current layer’s output would amount to $z^{\prime}_{j+1}=z^{\prime}_{j}+z_{j+1}$, but the sum of two Arcsin distributions turns into a badly behaved distribution (See Figure 2). Therefore, we introduce residual connections between the hidden vectors $h$, specifically as $h^{\prime}_{j}=(h_{j}+h^{\prime}_{j-1})/2$ with $z_{j+1}=A_{j}(h^{\prime}_{j})$.
Lemma B.1.
Let $X$,$Y\sim\mathcal{N}(0,1)$ be independent and $Z=(X+Y)/2$. Then $Z\sim\mathcal{N}(0,1)$.
Proof.
The sum of two independent normally distributed variables is equally normally distributed, with the mean and variance added up. This results in $\mathcal{N}(0+0,1+1)$. This is regularized to $\mathcal{N}(0,1)$ by dividing through the variance $2$. ∎
We assume independence between the two normally distributed vectors $h_{j}$ and $h^{\prime}_{j-1}$. With the residual sum being normally distributed again, we thus provide the requirements for Lemma 1.6. of Sitzmann et al. (2020) on keeping the activation distributions stable across the network.
Appendix C Training details
For each epoch, we iterate once over all land locations and sample a random month $m_{i}$ uniformly for each location. We train the network up to 20 epochs with the Adam optimizer and early stopping, totaling up to $547$ thousand training steps. We are learning a dense representation across the globe with a large number of datapoints and therefore do not require a val/test split. We use a model depth of 16, hidden size of 512, embeddings size of 256, and batch size of 8192 pixels. Although we found that a higher embedding size may slightly increase performance, we stick to 256 to align with similar geospatial encoding models. Our pretraining converges after $\sim$ 7 hours on a NVIDIA T4 with 16 GB of RAM and 16 CPUs.
Appendix D Downstream Tasks
Biomes classification.
We reproduce the biomes classification task from (Klemmer et al., 2023) by sampling 100k land locations uniformly across the globe and then extracting the biome at each location from the terrestrial ecoregions of the world map (Olson et al., 2001). The train-val-test split is $[0.5,0.1,0.4]$. As metric we use the mean of the F1 scores for all biomes, dubbed "macro" as it gives equal weight to all classes.
Species distribution modeling.
In species distribution modeling (SDM), we aim to predict if a species is likely to be observed at a location in a given month. We use the preprocessed (Dollinger et al., 2024) presence-only data from the GLC23 challenge (Botella et al., 2023), encompassing 2 million occurrences for 5.6 thousand plant species across Europe. The train-val-test split is $[0.7,0.05,0.25]$. We train the probe with the AN-full loss with background sampling (Cole et al., 2023), and report the Top-1 Accuracy.
Plant traits regression.
We sample 100k locations as in the biomes classification and extract means and standard deviations for 4 plant functional traits from the global plant trait maps provided by Moreno-Martínez et al. (2018). The traits are: specific leaf area, leaf dry matter content, leaf nitrogen and phosphorus content per dry mass, and leaf nitrogen/phosphorus ratio. We split the data $[0.5,0.1,0.4]$ and report the mean R² score across the eight values.
SDM has a date of observation associated with each occurrence, thus we sample Climplicit for the embedding of the respective month. In the case of biomes and plant traits, the rasters we sample from are immutable across the year. Here, we retrieve the embeddings for March, June, September and December from Climplicit and concatenate them as input to the probe for the classification and regression tasks.
Appendix E Future Work: Extension to Future Climatologies
CHELSA has been extended to include four monthly precipitation and temperature rasters for the climatologies 2011-2040, 2041-2070 and 2071-2100 (Karger et al., 2020). These are calculated from four different global circulation models for the three shared socioeconomic pathways SSP126, SSP370 and SSP585 (Gidden et al., 2019).
A future extension of Climplicit may include these future climatologies by first declaring a categorical variable $c_{i}\in\{-1,-0.33,0.33,1\}$ where each value corresponds to one of the four thirty-year intervals. This variable may be added to the positional encoding as another temporal variable:
$p_{i}=\Big{[}\>\lambda_{i},\>\phi_{i},\>\text{sin}(2\pi*m_{i}/12),\>\text{cos}(2\pi*m_{i}/12)\>,c_{i}\>\Big{]}\;\in\;[-1,1]^{5}\;.$
This may enable the location encoder to encode all four climatologies in a shared embedding space. It is important to note that for this method, not all eleven CHELSA variables from the contemporary climatology can be used during training, but only the four also provided for the future climatologies. This, along with the increased learning requirements from three additional climatologies might decrease performance on the downstream task setup of this work.
Having now learned a shared embedding space for the four climatologies, a downstream task may now train on the contemporary climatology $c_{i}=-1$, and make predictions on the future climatologies $c_{i}\in\{-0.33,0.33,1\}$. This can be seen as an downstream extrapolation to an un-sampled area, which the embeddings produced by location encoders are especially well suited to (Klemmer et al., 2023).
Appendix F Extended Results
Table 3: Linear probing on diverse geolocation representations for non-ecological downstream tasks. Mean and standard deviation of ten random initializations of the linear probe. Best. Best pretrained geolocation representation.
Model Pop Density (R² $\uparrow$) Med Income (R² $\uparrow$) Cali Housing (R² $\uparrow$)
FS Loc 75.3 $\pm$ 5.1 -7.5 $\pm$ 17.2 18.9 $\pm$ 18.9
FS CH 64.9 $\pm$ 2.1 29.8 $\pm$ 2.9 61.9 $\pm$ 2.2
FS Loc $+$ CH 75.1 $\pm$ 3.4 23.3 $\pm$ 13.5 60.3 $\pm$ 3.5
SatCLIP 49.1 $\pm$ 0.5 30.0 $\pm$ 9.1 35.2 $\pm$ 0.4
Taxabind 41.0 $\pm$ 0.2 15.0 $\pm$ 0.1 44.2 $\pm$ 0.1
SINR 38.8 $\pm$ 0.3 38.6 $\pm$ 0.5 37.1 $\pm$ 0.3
CSP 35.8 $\pm$ 0.7 27.3 $\pm$ 1.5 50.1 $\pm$ 0.6
GeoCLIP 39.3 $\pm$ 0.1 34.7 $\pm$ 0.1 58.9 $\pm$ 0.1
Climplicit 67.0 $\pm$ 0.3 45.0 $\pm$ 0.4 44.2 $\pm$ 0.6
F.1 Non-Ecological Downstream Tasks
We furthermore report in Table 3 the results on three non-ecological downstream tasks adapted from Klemmer et al. (2023). We sample 100k population density values in the United States of America from Doxsey-Whitfield et al. (2015). For median income regression, we use the 2021 US county-level median household income data from the US Department of Agriculture [^2] https://www.ers.usda.gov/data-products/county-level-data-sets/county-level-data-sets-download-data. California housing is a long-established dataset from Pace and Barry (1997) of median house prices in California census blocks. On all tasks we report the R2 score.
Climplicit is the best location encoder on two of the three tasks, also managing to outperform the fully trained models on the challenging median income task. Only on the California housing dataset does Climplicit fall short, which may be attributed to the high variation in housing prices in California independent of the local variation in climate.
F.2 Scaling of ReSIREN
Figure 3: Comparison of the scaling behavior of ReSIREN compared to SIREN. Generally, ReSIREN performs worse at few layers, but outperforms and scales better with more layers in exchange for an increase in training time.
In Figure 3 we report both the macro and weighted F1 score, along with the European SDM accuracy. Location and global CHELSA reconstruction are regression tasks where location have been sampled across the globe similar to the biomes task. For a location, either the location itself or the respective CHELSA vector was then regressed. Lastly, we present the increase in runtime from using the residual connections.
F.3 Qualitative Results
(a) Climplicit
(b) Groundtruth
(c) FS LOC
(d) FS CH
(e) FS LOC $+$ CH
(f) SatCLIP
(a) GeoCLIP
(b) CSP
(c) Taxabind
(d) SINR
Figure 5: Comparison of the fully-trained and pretrained embeddings by mapping the predicted biomes in the region around Lake Victoria, covering parts of Kenya, Tanzania, Uganda, Rwanda, Burundi and the Democratic Republic Kongo. Climplicit achieves a good reconstruction without any hallucinations, albeit missing high-level details such as the mangroves along the coast.
(a) March
(b) June
(c) September
(d) December
(e) FS LOC $+$ CH
(f) GeoCLIP
Figure 6: Visualization of the species distribution for Quercus ilex (Stone oak) for four months learned by probing the Climplicit embeddings in the SDM task. For each location, the model provides a probability of observing the oak (Blue: low, Yellow: high). The green dots represent the citizen scientist records of Quercus ilex for the respective month. Due to using months as input, the model learns different embeddings for each month, which allows for diverse distribution learning across the year. Additionally we present the distributions for the two strongest baselines.
(a) 2 layers
(b) 4 layers
(c) 8 layers
(d) 16 layers
(e) 32 layers
(f) SIREN 32 layers
Figure 7: Visualization of the species distribution for Quercus ilex (Stone oak) after probing the ReSIREN networks pretrained with different numbers of layers. All distributions are for the month of September. Blue to yellow signifies the predicted probability and green signifies the occurrences of the oak in September. It shows that with increasing depth, the level of detail also increases, though at the cost of increasing extrapolation (e.g. Skåne county in Sweden). We also present the distribution learned by a 32 layer SIREN.